OpenStudy (anonymous):

f(x) = 2x^4 − 4x^2 − 48 find the zeros

zepdrix (zepdrix):

To find the zeros, let f(x)=0, \[0=2(x^4-2x^2-24)\] Hmmmmm it looks like this might factor nicely :)

OpenStudy (anonymous):

darnit, too slow!! Follow @zepdrix .... :)

Directrix (directrix):

@yabucoayankee 2x^4 − 4x^2 − 48 = 0 Divide through by 2. x^4 - 2 x^2 - 24 = 0 Look for two numbers that multiply to -24 and sum to -2. Hint: If you can factor a^2 - 2a - 24, you can factor x^4 - 2 x^2 - 24